Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 16:09 Tuesday 20 August 2019 (10046) Print this report
Optic residual motion on Friday 16th of August.

On Friday 16th of August I measured the residual motion of the optic in ALIGNED state. There was a typhoon in the neighborhood and it was reported the BS was moving more than usual.

In the ALIGNED state the optic longitudinal RMS displacement and velocity DO NOT fulfill the requirement.

This table shows the integrated RMS displacement for each degree of freedom and the requirement.

Degree of freedom ALIGNED Requirement klog 9697
Longitudinal displacement 0.44 µm 0.4 µm 0.1 µm
Longitudinal velocity 0.8 µm/s 0.5 µm/s 0.14 µm/s
Pitch displacement 0.37 µrad 1 µrad 0.07 µrad
Yaw displacement 0.15 µrad 1 µrad 0.07 µrad

Most relevant notes:

  • Inertial damping on.
  • There's high coherence with the IP motion measured with the LVDTs from below 0.1 Hz up to 1.3 Hz. See plots attached.
  • The peak that appeared at 150-ish mHz the day before, appeared at 213-ish mHz. See entry 9991.

Other notes

  • Directory: /kagra/Dropbox/Subsystems/VIS/TypeBData/BS/Noise/Measurements/20190816/
  • File for ALIGNED state: BS_ALIGNED_TM_OPLEV_ASD_190816_1565941491.xml
  • Bandwidth of the displacement measurement' from 0.010 Hz to 800 Hz.
  • The bandwidth of the optic velocity measurement was from 0.010 Hz to 10 Hz. Above 10 Hz there is a lot of noise which I don't believe is real displacement of the optic. The peaks are likely related to the oplev optics support platform
  • Related entries: 9697.
Images attached to this report
Search Help